A mother and son are roller-skating together. The mother is standing still in the middle of the rink and the son runs into her from behind, skating as fast as he can. When he runs into his mother, he grabs onto her legs and they roll forward together at 2.07 m/s. The mother's mass is 61.7 kg, and the son has a mass of 30.1 kg. What is the son's velocity just prior to the collision?
1
Expert's answer
2019-01-24T10:23:07-0500

From the conservation of momentum:

mV=(m+M)v

Thus, the son's velocity just prior to the collision:


V=((m+M))/m v=(1+M/m)v=(1+61.7/30.1)2.07=6.31 m/s.